FEL Upgrade Project Weekly Brief

May 6-10, 2002




Special region drawings for the Recirculation Arcs, Wiggler, Injector dipoles, Optical Chicane and Light Box in FEL Upgrade were released to the Survey and Alignment Group this week with approved stand and cartridge positions so that stand installation work in these areas can proceed to completion.

We congratulate Chris Behre, a member of our Optics Group, who will be awarded his Baccalaureate Degree in Electrical engineering at Old Dominion University this weekend.


FEL Project Management was invited to make a presentation on FEL development to a DARPA Panel on the Electric Warship Design on May 14. G. Neil attended this week’s review of DOD’s Joint Technology Office funded programs at Albuquerque on May 6-8. We prepared abstracts for 6 presentations/posters that were submitted to the Program Committee today (May 10th) for the 2002 International FEL Conference that will be held in September 9-13 in Chicago. These abstracts will be forwarded to the ONR and AFRL program offices.

The AES/JLab CRADA to develop and test a 100 mA injector for high power FEL's has finally received full JLab approval and the Lab Director has submitted the CRADA to the DOE for approval.

Since CRADA numbers and dollars seems to be on the increase, the process by which such agreements are drafted, reviewed, and finalized needs to be re-examined to see if there are ways to streamline the process without sacrificing requirements. A proposal has been drafted and submitted to Tech Transfer Manager for comment and additional streamlining suggestions prior to submitting to members of the Directors Council.

WBS 3 (Beam Physics):

Another week of component review and analysis. This time, its the quads, which demonstrably can perform to spec but give different results for B(I) curves depending on the timing of the excitation (as expected). Unfortunately, we can't time the hysteresis loop in the same way in both the vault and on the test stand (different power supply systems), so the measured B(I) curve doesn't necessarily predict the achieved (in the vault)values to within spec (unfortunately, not as expected).

Discussions with MEG, CASA, controls, and EEG group members clarified the issue, and a plan to provide measurement capability matching operational practice is being worked up.

WBS 4 (Injector):

Continued design work on the new run/conditioning resistor. The support tube polishing was completed and is ready for implant. A step forward in the work with W&M on the implant process for the support tube was achieved. Reducing the operating pressure and a thorough cleaning of the plasma source chamber solved the problem of arcs observed in the first test tube. The second test tube shows no evidence of arcs after two runs. A third test run showed some arcing that was traced to dust that had accumulated at the interface of the appendage chamber holding the support tube when the chamber was flushed with gas. Since the gas can’t be flushed through the appendage, the flushing procedure will be replace with a chamber vacuum cleaning and a HEPA filter will be mounted above the chamber. A final test run with a new test tube will be done next week prior to commencing implantation on the actual cathode support tube. Three more runs are planned prior to implanting actual cathode support tube.

WBS 6 (RF):

Completed HV wiring of both CPSs for 100 kW klystrons with the exception of wiring to klystron carts.

The plumbing has been installed upstairs for the klystron carts on the LCW system. We have moved the 100 kW klystron cart into position for unit #3 due to the relocation of the magnet box power supply into FL01 zone.

We have received the gun tank chassis support struts.


WBS 8 (Instrumentation):

1) Cable tray straightening was completed

2) 1st article 8 channel stepper motor chassis was turned over to Al for system debug. (good results)

3) IOC FEL3 was upgraded to Epics r3132j0 and the timer card software and hardware was changed out to provide a one-board system. Preliminary testing is good.

4) CANbus interface for the beamviewers was turned over to Jeff Wilson with a completion date around May 30.

5) Electrical plumbing for the new dipole box supply has begun.

6) The rest of the 3" beam viewers have had the interface box converted, tested and turned over to the vacuum group(c/o J. Heckman) for installation.

7) The water supply (feed/return, manifolds and valves) for the super trim magnet racks have been identified and some parts ordered.

8) The prexisting beamviewer cameras CCD's have been tested and repaired.

9) Re-terminated trim magnet feeds in Junction Box 1and 2.

10) Fabrication of four 4 channel programmable digital pulse generator boards were completed





WBS 9 (Transport):


Optical Chicane Dipoles (GW)

• At Magnet Enterprises International (MEI) in Oakland, CA. We examined the first eight coils. They are electrically ok but have some surface imperfections that require repair. They have potted their ninth coil with some changes in method that should avoid the imperfections in future coils.

• Upper and lower core’s plates are in final machining with projection that first article assembly will be in a week and a half.

• They have made the fixtures for gluing the Purcell gap shims in place and they have some trial shim pieces and a trial core piece that they will do some trials with before they will do the real gluing.

Injector Dipoles (DU/DV)

• Core pieces are in final machining at Master Machine.

• All coils are wound and potting fixture

Arc Dipoles (GY, GX, GQ)

• Process Equipment Co. continued manufacturing the cores and parts of the GQ-GX dipoles. We will visit them next week to go over their progress.

• Wang NMR continues working on the winding fixture for the coils.

• AES completed the run GX dipoles at high energy with a larger shim only on the IR portion of the large face. Dave Douglas is evaluating it. They have started drawing Rev A to the drawings that incorporates the shims and their holding mechanisms. This has no effect on present production because the modifications only apply at final machining.

• Wang NMR completed winding the first GY coil. With this experience, they expect the next to take three weeks and the final two to take two weeks each.

• Bosma Machine will complete the first article GY Core on Monday. I will visit them on Wednesday to approve.


3 inch quad (QX)

• We decided to re measure all the QX quads to the new CEBAF, full range ("bang-bang") hysteresis protocol. However, further tests showed that the power supply we use and the trim rack power supplies do not have the same characteristics. At weeks end we are considering our options.

Trim Quad (QT)

• Milhous Control of Virginia continues making the remaining seven magnets. We found that the first article had a leak in the cooling circuit from a post manufacturing clean out of a water fitting.

• As a quick analog computation, (Using the device itself) we tested the quad at low current to assess whether by powering coils appropriately it would be able to have the superposed vertical corrector GC built into it. Results were such that we will try it on our magnetic model.

Sextupole (SF)

• DULY Research continues to work on completing drawings of the magnet.

Octupole (OT)

• Work on this magnet is on furlough until the sextupole is designed.

Beam Line and Vacuum

• The drawings for ARC Chambers have been checked and are being revised to reflect the comments.

• Special region drawings for Arcs, Wiggler, Injector dipoles, Optical Chicane and Light Box were given to the Survey and Alignment Group with preliminary graphics but approved stand and cartridge positions following a final check.

• Additional personnel were assigned to the recirculation region in front of the Cryomodule.

• The optical chicane region with its chamber continues in development.

WBS 11 (Optics):

We met with members of the Survey and Alignment group to determine if the number and size of access ports in the optical cavity vacuum vessels is adequate. Agreement was reached on a strawman alignment plan, and this is incorporated into the drawings. An assembly and test fixture was partially "mocked up" in User Lab 2, and based on the layout, sketches created and submitted to the machine shop for fabrication. The electrical feedthroughs for the vacuum vessels have been ordered. We received the leadscrews and nuts for the optical cavities and the assembly/test stands.

We have a scheme for translating the flexible cooling water hoses and electrical conduit, and have ordered parts to test it. We are designing the backplane cooling for the mirror and turning cassettes. We have a budgetary quote for the insertable mirrors (which are water cooled) and the bid seems high so we are checking other sources.

While we achieve acceptable performance from the feedback loop on the OCMMS (active stabilization) electronics, we would like it to be ~ 5X better. Plans for the new circuit were made; smaller and higher resolution position sensitive detectors were ordered. Work continues on the drive laser pulse compressor. We installed a spectrometer in the drive laser room to measure the bandwidth, and thus to determine the pulsewidth. Sketches to provide mounting fixtures to In-braze mirrors and work with Ga:In eutectic were submitted to the machine shop for fabrication. The vendor of our ultrafast laser system informed us that the Ti:sapphire oscillator should ship today.

We attended meetings reviewed plans for laser purchases, and created a preliminary design for an external resonant cavity for the FEL (aka the pulse stacking cavity). This is primarily funded by a DIRUP grant to the College of William and Mary (M. Kelley - P.I.)

We prepared several abstracts of papers to present at the FEL conference this September.

Our active mirror stabilization effort has involved a great deal of work by an Old Dominion Univ. EE student, Mr. Chris Behre. Chris prepared and gave a presentation before a panel of three engineers, and prepared a paper as well. His work was awarded a 2nd place out of 17 entries. We're all proud of Chris, and congratulate him on his graduation, which takes place this weekend.